# Piecewise Smooth Dynamical Systems Regularized by Convolution

### v1

@@ 1 Introduction
In a series of papers, [16, 17, 18, 19], Llibre Silva and Teixeira studied several regularization problems in the $\mathbbm{R}^{n}$, including the double regularization in the case in which the discontinuity has codimension 2 (intersection of two planes) and the regularization in more degenerate surfaces.
@@
Let us briefly explain the this result. Firstly, we observe that the regularized family $X_{\varepsilon}$ can be seen as a vector field in the product space $N=M\times(\mathbbm{R}_{\geq 0},0)$, with discontinuity locus $\Sigma\times\{0\}$. Then, performing a single blowing-up on $N$,
@@
The regularization by convolution is a very classical tool from analysis, which can be defined in a much broader context. Let us briefly describe the construction and refer to subsection 3.1 for more details: Let $X$ be a locally integrable vector field in $\mathbbm{R}^{n}$. In other words, we assume that $X$ has the form
@@
such that the pull-back of $X^{\mathrm{reg}}$ under the composition $\Phi=\varphi_{r}\circ\cdots\circ\varphi_{1}$ extends to a smooth oriented 1-foliation on $\widetilde{N}$. For this reason, we refer to such sequence of blowing-ups as a smoothing procedure for the regularization $X^{\mathrm{reg}}$ of $X$.
@@ 1.1 Overview of the paper
which generates a one-parameter family of smooth functions depending on $\varepsilon>0$ which we complete to $\varepsilon=0$ by defining $f_{0}=f$. We show that this process can be interpreted as a smooth extension in the augmented space $N=M\times\mathbb{R}_{\geq 0}$, a manifold with corners whose boundary component $M=\{\varepsilon=0\}$ represents the initial discontinuous domain.
@@ 4.3.1 Phase-directional blowing-up
where $\underline{s}\,(x_{I}-t_{I}\varepsilon)>0$ stands for the subset of $\mathbbm{R}^{c}$ defined by the conditions $s_{1}(x_{1}-t_{1}\varepsilon)>0,\ldots,s_{c}(x_{c}-t_{c}\varepsilon)>0)$.
On the other hand, the directional blowing-up $B_{(I,i_{1})}$ has the expression

### v2

@@ 1 Introduction
In a series of papers, [17, 18, 19, 20], Llibre, Silva and Teixeira studied several regularization problems in the $\mathbbm{R}^{n}$, including the double regularization in the case where the discontinuity set is given by the union of two transversal hyperplanes and the regularization in more degenerate surfaces.
@@
Let us briefly explain this result. Firstly, we observe that the regularized family $X_{\varepsilon}$ can be seen as a vector field in the product space $N=M\times(\mathbbm{R}_{\geq 0},0)$, with discontinuity locus $\Sigma\times\{0\}$. Then, performing a single blowing-up on $N$,
@@
The regularization by convolution is a very classical tool from analysis, which can be defined in a much broader context. We now briefly describe the construction and refer to subsection 3.1 for more details: Let $X$ be a locally integrable vector field in $\mathbbm{R}^{n}$. In other words, we assume that $X$ has the form
@@
such that the pull-back of $X^{\mathrm{reg}}$ under the composition $\Phi=\varphi_{r}\circ\cdots\circ\varphi_{1}$ extends to a smooth oriented 1-foliation on $\widetilde{N}$, which is a manifold with corners. For this reason, we refer to such sequence of blowing-ups as a smoothing procedure for the regularization $X^{\mathrm{reg}}$ of $X$.
@@ 1.1 Overview of the paper
which generates a one-parameter family of smooth functions depending on $\varepsilon>0$ which we complete to $\varepsilon=0$ by defining $f_{0}=f$. We show that this process can be interpreted as a smooth extension in the augmented space $N=M\times\mathbb{R}_{\geq 0}$, a manifold with boundary whose boundary component $M=\{\varepsilon=0\}$ represents the initial discontinuous domain.
@@ 3 Regularization by convolution
This section introduces the main technical tools used in the paper. We begin by recalling the basic definitions of manifolds with corners, directional blow-up maps, and regularization by convolution. The main result is the smoothing procedure, which is proved in Theorem 4.12.
@@ 3.2 Regularization of piecewise-smooth functions
Remark 3.4.
(1) Note that the convolution integral $f_{\varepsilon}=m_{\varepsilon}\ast f$ is written in terms of the global Euclidean coordinates of $\mathbb{R}^{n}$. For our purposes, it will be necessary to express this integral in local adapted coordinates, as in (8), which are defined in the vicinity of each point in discontinuity locus.
More precisely, let $(U,\varphi)$ be a local chart defining adapted coordinates, and let $\psi=\varphi^{-1}$ denote the inverse diffeomorphism, with domain on the open subset $V=\varphi(U)$ of $\mathbbm{R}^{n}$. The goal is to write the pull-back of the regularized function $f_{\varepsilon}\circ\psi$ in terms of a convolution-type integral involving the pull-back of the piecewise smooth function $f\circ\psi$.
By a change of variables in the convolution integral (11) we can write
$(f_{\varepsilon}\circ\psi)(x)=\int_{\mathbb{R}^{n}}(f\circ\psi)(x-t)\,\boldsymbol{m}_{\varepsilon,x}(t)\,dt.$ (12)
Here $t\mapsto\boldsymbol{m}_{\varepsilon,x}(t)$ is a family of functions depending smoothly on $\varepsilon,x$, with domain some neighborhood of the origin in $\mathbb{R}^{n}$. More precisely, we can write
$\boldsymbol{m}_{\varepsilon,x}(t)=\frac{1}{\varepsilon^{n}}\,u(x,t)\,m\!\left(\frac{t\,v(x,t)}{\varepsilon}\right),$
where $m$ is the original mollifier appearing in (10) and $u,v$ are smooth functions on some open neighborhood of the origin in $V\times\mathbb{R}^{n}$ that do not vanish at $(x,t)=(0,0)$.
Note that, for each fixed value of $x$, the support of the function $t\mapsto\mathbf{m}_{\varepsilon,x}$ shrinks to the origin as $\varepsilon\to 0$. Consequently, for $\varepsilon>0$ sufficiently small, the above integral is well defined for all $t\in\mathbb{R}^{n}$.
In other words, when expressed in the local adapted coordinates $x$, the regularized function $f_{\varepsilon}\circ\psi$ can be written as the convolution of $f\circ\psi$ with a generalized mollifier $\boldsymbol{m}_{\varepsilon,x}$, which, in these coordinates, depends both on the integration variable $t\in\mathbb{R}^{n}$ and on the base point $x$.
For the computations that follow, this additional dependence of the mollifier $\boldsymbol{m}_{\varepsilon,x}$ on the base point is entirely harmless. Accordingly, we shall omit it from the notation.
(2) The regularization by convolution can be defined more generally for an arbitrary Riemannian manifold $M$ through the Greene-Wu convolution formula
$f_{\varepsilon}(p)=\int_{v\in T_{p}M}f(\exp_{p}v)\,m_{\varepsilon}(v)dv$
where $\exp_{p}:T_{p}M\rightarrow M$ denotes the exponential map associated to the metric and the integral is computed with respect to the Lebesgue measure on $T_{p}M$ induced by the metric. In this more general setting, the Greene–Wu convolution formula expressed in local charts has the same form as (12). We refer the reader to [13] for further details.
@@ 4.3.1 Phase-directional blowing-up
where the notation $\{\underline{s}\,(x_{I}-t_{I}\varepsilon)>0\}$ stands for the subset of $\mathbbm{R}^{c}$ defined by the conditions $s_{1}(x_{1}-t_{1}\varepsilon)>0,\ldots,s_{c}(x_{c}-t_{c}\varepsilon)>0$. We recall that the convention established in Remark 3.4 is in force for all subsequent computations.
Observe that the directional blowing-up $B_{(I,i_{1})}$ has the expression